# Semantic MIMO Systems for Speech-to-Text Transmission

### v1

@@ III-A SU-MIMO Communicaton System
III-A1 First training stage
The goal of the first training stage is to train the DeepSC-ST 2.0 that recovers the text sequence at the receiver from the input speech over MIMO channels with SVD precoding. As shown in Fig. 3(a), the transformer-enabled semantic encoder converts the input speech, $\boldsymbol{s}$, into the low-dimensional semantic features, $\boldsymbol{f}$. The channel encoder processes $\boldsymbol{f}$ and produces the symbols, $\boldsymbol{x}$, that are precoded into $\widetilde{\boldsymbol{x}}$ to be transmitted over the MIMO channel, $\boldsymbol{H}$. Therefore, $\widetilde{\boldsymbol{x}}$ can be denoted as
$\widetilde{\boldsymbol{x}}$ is transmitted to the user over the MIMO channel according to (2). As aforementioned that the SVD precoding-based MIMO channel can be regarded as multiple SISO channels. The received symbols, $\widetilde{\boldsymbol{y}}$, are passing through the channel decoder to obtain the recovered semantic features, $\widehat{\boldsymbol{f}}$. Finally, the semantic decoder takes $\widehat{\boldsymbol{f}}$ as the input and produces the text sequence, $\widehat{\boldsymbol{t}}$, to serve the users only requesting the text information. The process to obtain $\widehat{\boldsymbol{t}}$ from $\widetilde{\boldsymbol{y}}$ can be expressed as
@@ III-A1 First training stage
where $\mathfrak{A}(\boldsymbol{s},\boldsymbol{\;}\boldsymbol{t})$ is all the alignments from $\boldsymbol{t}$ to $\boldsymbol{s}$ and $\boldsymbol{t}$ is the correct text sequence.
@@ III-A2 Second training stage
A semantic-aware network is developed in the second training stage to analyse the semantic features and return their importance. Particularly, the trained semantic encoder from the first training stage is leveraged to process input speech $\boldsymbol{s}$ and produce semantic features $\boldsymbol{f}$. Then, $\boldsymbol{f}$ is fed into the semantic-aware network to learn the importance matrix, $\boldsymbol{w}$, that consists of the importance coefficients of corresponding semantic features. According to the importance matrix $\boldsymbol{w}$, the semantic features containing the critical information in $\boldsymbol{f}$ are preserved while other semantic features are masked to 0, which generates the selected semantic features, $\overline{\boldsymbol{f}}$. The trainable parameters of the semantic-aware network is ${\boldsymbol{\theta}}_{2}$, then $\overline{\boldsymbol{f}}$ is produced by
@@
The motivations behind the loss function ${\mathcal{L}}_{\mathrm{SA}}$ is to ensure that the selected semantic features, $\overline{\boldsymbol{f}}$, simultaneously minimize the information loss compared to semantic features $\boldsymbol{f}$ and predict the accurate text sequence as much as possible. By doing so, the essential semantic features are identified based on the importance matrix, $\boldsymbol{w}$, and these features occupy high priority to be transmitted over good SISO channels. The training algorithm of the semantic-aware network is described in Algorithm 2. Note that the trainable parameters of the semantic encoder are not updated during this training stage.
III-A3 Testing Stage
In the testing stage, the trained ${\mathfrak{T}}_{\mathrm{S}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{C}}(\cdot)$, $\mathfrak{T}_{\mathrm{C}}^{-1}(\cdot)$, and $\mathfrak{T}_{\mathrm{S}}^{-1}(\cdot)$ from the first training stage, as well as the trained ${\mathfrak{T}}_{\mathrm{SA}}(\cdot)$ from the second stage, are used to facilitate the semantic transmission over MIMO channels. Particularly, the semantic features, $\boldsymbol{f}$, are produced by the semantic encoder and fed into the semantic-aware network to obtain the importance matrix, $\boldsymbol{w}$, which enables to sort $\boldsymbol{f}$ according to $\boldsymbol{w}$ and to generate the sorted semantic features, $\vec{\boldsymbol{f}}$, in the descending order. Based on $\vec{\boldsymbol{f}}$, the encoded symbols in $\vec{\boldsymbol{x}}$ that correspond to the critical semantic information are transmitted over SISO channels with high SNR values, which minimizes the semantic error and improves the intelligibility of the recovered text sequence, $\widehat{\boldsymbol{t}}$. The testing algorithm of the proposed SAC-ST in the SU-MIMO communication system is described in Algorithm 3. According to Algorithm 3, it is worth mentioning that the recovered semantic features, $\widehat{\vec{\boldsymbol{f}}}$, are resorted to the original order before passing through the semantic decoder. The symbols $\widetilde{\vec{\boldsymbol{x}}}\in\mathfrak{R}^{B\times D}$, where $B$ is the number of semantic features, are converted to complex values and reshaped to $\mathfrak{C}^{\left(B/N\right)\times N\times\left(D/2\right)}$ prior to transmission. In each time slot, the symbols transmitted over each SISO channel correspond to one semantic feature, denoted as ${\widetilde{\vec{\boldsymbol{x}}}}_{b}\in\mathfrak{C}^{1\times\left(D/2\right)}$.
@@ III-A3 Testing Stage
4:     $sf({\mathfrak{T}}_{\mathrm{SA}}(\boldsymbol{f});\kappa)\rightarrow\overline{\boldsymbol{f}}$.
@@
9:              Transmit ${\widetilde{\vec{\boldsymbol{x}}}}_{b}$ over $j$-th SISO channel via (2)
@@ III-B MU-MIMO Communicaton System
Based on the DeepSC-ST 2.0 and the fine-tuned semantic-aware network, the proposed SAC-ST achieves speech-to-text transmission over the MU-MIMO channel, which satisfies the requirements of different users requesting text information simultaneously. As aforementioned, the information of other users experiences the MUI and channel impairment when the essential information of $k$-th user is recovered accurately. In other words, a high semantic fidelity transmission for one certain user can be ensured by transmitting all the critical semantic features of this user over good SISO channels. However, it diminishes the priority of other users to recover the essential information, which runs count to the ultimate goal of a satisfactory quality of experience for all users. To maintain fairness, each user occupies an equal probability of transmitting their semantic features over good SISO channels. The testing process of SAC-ST in the MU-MIMO communication system is described in Algorithm 4.
@@
12:              Transmit ${\widetilde{\vec{\boldsymbol{x}}}}_{k,b1}$ over $j$-th SISO channel
@@ III-C Channel Estimation Network
To render the proposed SAC-ST applicable to practical communication environments, our endeavours extend to the estimation of $\boldsymbol{H}$ in the SU-MIMO and MU-MIMO communication systems by leveraging a neural network-enabled channel estimation network, named ChanEst network. This network is responsible for providing an approximate channel, $\widehat{\boldsymbol{H}}$, for the seamless integration of the precoding process in the BS before transmission and breaking the dependence on perfect CSI, as shown in Fig. 5. From the figure, the pilot symbols, ${\boldsymbol{x}}_{p}$, are first transmitted from receiver users to the BS over the real MIMO channel $\boldsymbol{H}$. The received symbols, ${\boldsymbol{y}}_{p}$, are processed by the conventional channel estimation approach to obtain the pre-estimated MIMO channel, $\widetilde{\boldsymbol{H}}$. The ChanEst network takes $\widetilde{\boldsymbol{H}}$ as input and generates the final estimated channel $\widehat{\boldsymbol{H}}$. Then the symbols, $\boldsymbol{x}$, are precoded to attain $\widetilde{\boldsymbol{x}}$ according to $\widehat{\boldsymbol{H}}$. $\widetilde{\boldsymbol{x}}$ is transmitted over $\widehat{\boldsymbol{H}}$ and the received symbols is $\widetilde{\boldsymbol{y}}$. The mean-squared error (MSE) is adopted as the loss function to train the ChanEst network, expressed as
@@ IV Numerical Results
In this section, we test the performance of the proposed SAC-ST over SU-MIMO and MU-MIMO Rayleigh fading channels, where the word error-rate (WER) and sentence similarity [38] are utilized as performance metrics. We compare SAC-ST to the DeepSC-ST [18] and the framework without the semantic-aware network, i.e., DeepSC-ST 2.0. Moreover, the simulation results of the SAC-ST with the ChanEst network are presented. LibriSpeech corpus is adopted in the experiments, and the simulation environment is Pytorch 1.13.1 and NVIDIA GPU A100.
@@
IV-A Neural Network Settings
As aforementioned that the DeepSC-ST 2.0 and the semantic-aware network are trained independently. In the DeepSC-ST 2.0, 12 transformer encoder layers are concatenated to construct the semantic encoder. Three dense layers with 512 units are utilized in both the channel encoder and the channel decoder. The semantic decoder consists of a dense layer containing 32 units followed by a softmax activation function. In the semantic-aware network, three dense layers with 32 units are leveraged. The ChanEst network includes three 2D CNN layers containing 64 units, and the kernel size of each layer is $5\times 5$. The batch size is 64, and the number of warm-up steps is 3000 for training all the neural networks. The parameter settings of the SAC-ST and the ChanEst network are summarized in Table I.
@@ IV-A Neural Network Settings
SAC-ST DeepSC-ST 2.0 Semantic Encoder 12×12$\times$12 ×Transformer Encoder 128 (8 heads) GELU
Channel Encoder 2×2$\times$2 ×Dense 512 units ReLU6
Dense 512 units None
Channel Decoder 3×3$\times$3 ×Dense 512 units ReLU6
Semantic Decoder Dense 32 units Softmax
Semanitc-Aware Network 2×2$\times$2 ×Dense 32 units ReLU6
Dense 32 units None
ChanEst Network 2×2$\times$2 ×2D CNN 64 units HardTanh
2D CNN 64 units None
@@ IV-B SU-MIMO Experiments
The WER results of the proposed SAC-ST in the SU-MIMO transmission system are shown in Fig. 6(a), where perfect CSI is assumed. The ground truth is the WER score obtained by feeding the speech input into the speech-to-text pipeline, Conformer [39], directly and no speech transmission problems are considered. From the figure, the upgraded DeepSC-ST 2.0 achieves lower WER scores than DeepSC-ST due to the improvement of the attention mechanism by incorporating the transformer module. The proposed SAC-ST outperforms the DeepSC-ST 2.0 and obtains a decrease of around $75.5\%$ when SNR$=$-8 dB. This is attributed to the semantic-aware network by transmitting the essential semantic information over good SISO channels. Besides, the minimum SNR value to enable the readable sentence in the SAC-ST is around -8 dB, while it is around 4 dB in the DeepSC-ST 2.0. Moreover, the gap between the WER scores of the DeepSC-ST 2.0 and the proposed SAC-ST gradually narrows when the SNR value increases because the important semantic features in $\boldsymbol{f}$ are recovered accurately in both schemes when SNR$>$13 dB.
@@
Fig. 6(b) illustrates the sentence similarity scores of different approaches. From the figure, the sentence similarity score of the DeepSC-ST and the DeepSC-ST 2.0 is nearly 0.2 for SNR$=$-8 dB, while the SAC-ST attains a score of 0.6. Similar to the result of WER, the proposed SAC-ST significantly outperforms the DeepSC-ST 2.0 in the low SNR regime, and the difference between them decreases when SNR$>$13 dB, which proves the effectiveness of the semantic-aware network in improving the semantic fidelity by perceiving the semantic importance over MIMO transmission paradigm with SVD precoding. The representative results of the DeepSC-ST, the DeepSC-ST 2.0, and the SAC-ST are shown in Table II, where the error words are highlighted in red. From the table, the sentence of the DeepSC-ST is totally corrupted and unable to read, and limited information is obtained from a few correct words scattered sporadically throughout the sentence of the DeepSC-ST 2.0. However, the sentence of the proposed SAC-ST is very intelligible with high semantic fidelity.
@@
Next, we investigate the antenna diversity of SAC-ST to cope with various antenna distributions in realistic SU-MIMO communication scenarios, and the results of WER scores and sentence similarity scores are shown in Fig. 7. From the figure, SAC-ST has superior performance and is more robust to channel impairments in the transmission paradigm with more antennas. To verify that, we calculate the average equivalent SNR values, $\widetilde{\boldsymbol{\sigma}}$, in all SISO channels for different antenna combinations, and the example when true SNR value $\sigma$ is -8 dB is shown in Table III [^1] All the data of the equivalent SNR values $\widetilde{\boldsymbol{\sigma}}$ for different antenna combinations and for true SNR value $\sigma$ from -8 dB to 22 dB is available at https://github.com/Zhenzi-Weng/SAC-ST/tree/main/equivalent-SNR.. From the Table, the SNR values in the top SISO channels of the antenna combination of $16\times 16$ are much higher than that of $2\times 2$, which facilitates more accurate semantic fidelity transmission. Moreover, as the number of antennas increases, the performance improvement between two different antenna combinations becomes negligible due to the insignificant difference between their SNR values in the top SISO channels.
TABLE II: Recovered sentences of the DeepSC-ST, the DeepSC-ST 2.0, and the SAC-ST when SNR is -2 dB.
@@
TABLE III: Equivalent SNR values $\widetilde{\boldsymbol{\sigma}}$ for different antenna combinations when true SNR value $\sigma$ is -8 dB.
@@
$N\times M$ Equivalent SNR values $\widetilde{\boldsymbol{\sigma}}$ in SISO channels (dB)
@@
The experimental results of the SAC-ST with different channel estimation algorithms are shown in Fig. 8. We consider three benchmarks. The first and second benchmarks test the SAC-ST trained under perfect CSI in the SU-MIMO transmission system with the conventional least square (LS) and minimum mean-square error (MMSE) channel estimation algorithms, respectively. The third benchmark is obtained by training and testing the SAC-ST under the channel environments estimated through the MMSE algorithm. From the figure, the SAC-ST with the ChanEst network outperforms all the benchmarks, especially in the low SNR regime, and it attains the WER scores and the sentence similarity scores comparable to the SAC-ST with perfect CSI. The representative results of the SAC-ST with various channel estimation methods in the SU-MIMO transmission system are shown in Table IV.
(a) WER
(b) Sentence similarity
TABLE IV: Recovered sentences of the SAC-ST with different channel estimation algorithms when SNR is -8 dB.

### v2

@@ III-A1 First training stage
$\widetilde{\boldsymbol{x}}$ is transmitted to the user over the MIMO channel according to (2). As aforementioned that the SVD precoding-based MIMO channel can be regarded as multiple SISO subchannels. The received symbols, $\widetilde{\boldsymbol{y}}$, are passing through the channel decoder to obtain the recovered semantic features, $\widehat{\boldsymbol{f}}$. Finally, the semantic decoder takes $\widehat{\boldsymbol{f}}$ as the input and produces the text sequence, $\widehat{\boldsymbol{t}}$, to serve the users only requesting the text information. The process to obtain $\widehat{\boldsymbol{t}}$ from $\widetilde{\boldsymbol{y}}$ can be expressed as
@@
where $\mathfrak{A}(\boldsymbol{s},\boldsymbol{\;}\boldsymbol{t})$ is all the alignments from $\boldsymbol{t}$ to $\boldsymbol{s}$ and $\boldsymbol{t}$ is the correct text sequence. Assume $\boldsymbol{t}=\left[\mathrm{t},\;\mathrm{h},\;\mathrm{i},\;\mathrm{s}\right]$, the valid alignments involve $\left[blank,\;\mathrm{t},\;blank,\;\mathrm{h},\;\mathrm{i},\;\mathrm{s},\;blank\right]$, or $\left[\mathrm{t},\;blank,\;\mathrm{h},\;\mathrm{i},\;blank,\;blank,\;\mathrm{s}\right]$, etc, because the $blank$ token is eliminated when attaining $\widehat{\boldsymbol{t}}$.
@@ III-A2 Second training stage
A semantic-aware network is developed in the second training stage to analyse the semantic features and return their importance. Particularly, the trained semantic encoder from the first training stage is leveraged to process input speech $\boldsymbol{s}$ and produce semantic features $\boldsymbol{f}$. Then, $\boldsymbol{f}$ is fed into the semantic-aware network to learn the importance matrix, $\boldsymbol{w}$, that consists of the importance coefficients of corresponding semantic features. According to $\boldsymbol{w}$, the features in $\boldsymbol{f}$ corresponding to high importance coefficients in $\boldsymbol{w}$ are preserved, while the features in $\boldsymbol{f}$ corresponding to low importance coefficients in $\boldsymbol{w}$ are masked to zero, which returns the selected semantic feature, $\overline{\boldsymbol{f}}$. Note that the order of semantic features in $\overline{\boldsymbol{f}}$ is consistent with $\boldsymbol{f}$. The trainable parameters of the semantic-aware network is ${\boldsymbol{\theta}}_{2}$, then $\overline{\boldsymbol{f}}$ is produced by
@@
The motivations behind the loss function ${\mathcal{L}}_{\mathrm{SA}}$ is to ensure that the selected semantic features, $\overline{\boldsymbol{f}}$, simultaneously minimize the information loss compared to semantic features $\boldsymbol{f}$ and predict the accurate text sequence as much as possible. By doing so, the essential semantic features are identified based on the importance matrix, $\boldsymbol{w}$, and these features occupy high priority to be transmitted over good SISO subchannels. The training algorithm of the semantic-aware network is described in Algorithm 2. Note that the trainable parameters of the semantic encoder are not updated during this training stage.
III-A3 Testing Stage
In the testing stage, the trained ${\mathfrak{T}}_{\mathrm{S}}(\cdot)$, ${\mathfrak{T}}_{\mathrm{C}}(\cdot)$, $\mathfrak{T}_{\mathrm{C}}^{-1}(\cdot)$, and $\mathfrak{T}_{\mathrm{S}}^{-1}(\cdot)$ from the first training stage, as well as the trained ${\mathfrak{T}}_{\mathrm{SA}}(\cdot)$ from the second stage, are used to facilitate the semantic transmission over MIMO channels. Particularly, the semantic features, $\boldsymbol{f}$, are produced by the semantic encoder and fed into the semantic-aware network to obtain the importance matrix, $\boldsymbol{w}$. To enable the essential semantic features in $\boldsymbol{f}$ to be transmitted over top subchannels, $\boldsymbol{f}$ is permuted in descending order based on their corresponding importance coefficients in $\boldsymbol{w}$, which returns the sorted semantic features, $\vec{\boldsymbol{f}}$. The order of semantic features in $\vec{\boldsymbol{f}}$ is inconsistent with $\boldsymbol{f}$. $\vec{\boldsymbol{f}}$ is fed into the channel encoder to attain the symbols, $\vec{\boldsymbol{x}}$. According to $\vec{\boldsymbol{f}}$, the encoded symbols in $\vec{\boldsymbol{x}}$ that correspond to the critical semantic information are transmitted over SISO subchannels with high SNR values, which minimizes the semantic error and improves the intelligibility of the recovered text sequence, $\widehat{\boldsymbol{t}}$. The testing algorithm of the proposed SAC-ST in the SU-MIMO communication system is described in Algorithm 3. According to Algorithm 3, it is worth mentioning that the recovered semantic features, $\widehat{\vec{\boldsymbol{f}}}$, are resorted to the original order before passing through the semantic decoder. The symbols $\widetilde{\vec{\boldsymbol{x}}}\in\mathfrak{R}^{B\times D}$, where $B$ is the number of semantic features, are converted to complex values and reshaped to $\mathfrak{C}^{\left(B/N\right)\times N\times\left(D/2\right)}$ prior to transmission. In each time slot, the symbols transmitted over each SISO subchannel correspond to one semantic feature, denoted as ${\widetilde{\vec{\boldsymbol{x}}}}_{b}\in\mathfrak{C}^{1\times\left(D/2\right)}$.
@@ III-A3 Testing Stage
4:     $sf({\mathfrak{T}}_{\mathrm{SA}}(\boldsymbol{f});\mu)\rightarrow\overline{\boldsymbol{f}}$.
@@
9:              Transmit ${\widetilde{\vec{\boldsymbol{x}}}}_{b}$ over $j$-th SISO subchannel via (2)
@@ III-B MU-MIMO Communicaton System
12:              Transmit ${\widetilde{\vec{\boldsymbol{x}}}}_{k,b1}$ over $j$-th SISO subchannel
@@
Based on the DeepSC-ST 2.0 and the fine-tuned semantic-aware network, the proposed SAC-ST achieves speech-to-text transmission over the MU-MIMO channel, which satisfies the requirements of different users requesting text information simultaneously. As aforementioned, the information of other users experiences the MUI and channel impairment when the essential information of $k$-th user is recovered accurately. In other words, a high semantic fidelity transmission for one certain user can be ensured by transmitting all the critical semantic features of this user over good SISO subchannels. However, it diminishes the priority of other users to recover the essential information, which runs count to the ultimate goal of a satisfactory quality of experience for all users. To maintain fairness, each user occupies an equal probability of transmitting their semantic features over good SISO subchannels. The testing process of SAC-ST in the MU-MIMO communication system is described in Algorithm 4.
@@ III-C Channel Estimation Network
The DL techniques have shown great advantages compared to conventional channel estimation algorithms [43, 44, 45]. To render the proposed SAC-ST applicable to practical communication environments, our endeavours extend to the estimation of $\boldsymbol{H}$ in the SU-MIMO and MU-MIMO communication systems by leveraging a neural network-enabled channel estimation network, named ChanEst network. This network is responsible for providing an approximate channel, $\widehat{\boldsymbol{H}}$, for the seamless integration of the precoding process in the BS before transmission and breaking the dependence on perfect CSI, as shown in Fig. 6. From the figure, the pilot symbols, ${\boldsymbol{x}}_{p}$, are first transmitted from receiver users to the BS over the real MIMO channel $\boldsymbol{H}$. The received symbols, ${\boldsymbol{y}}_{p}$, are processed by the conventional channel estimation approach to obtain the pre-estimated MIMO channel, $\widetilde{\boldsymbol{H}}$. The ChanEst network takes $\widetilde{\boldsymbol{H}}$ as input and generates the final estimated channel $\widehat{\boldsymbol{H}}$. Then the symbols, $\boldsymbol{x}$, are precoded to attain $\widetilde{\boldsymbol{x}}$ according to $\widehat{\boldsymbol{H}}$. $\widetilde{\boldsymbol{x}}$ is transmitted over $\widehat{\boldsymbol{H}}$ and the received symbols is $\widetilde{\boldsymbol{y}}$. The mean-squared error (MSE) is adopted as the loss function to train the ChanEst network, expressed as
@@ IV Numerical Results
In this section, we test the performance of the proposed SAC-ST over SU-MIMO and MU-MIMO Rayleigh fading channels, where the word error-rate (WER) and sentence similarity [46] are utilized as performance metrics. We compare SAC-ST to the DeepSC-ST [20] and the framework without the semantic-aware network, i.e., DeepSC-ST 2.0. Moreover, the simulation results of the SAC-ST with the ChanEst network are presented. LibriSpeech 960h corpus is adopted in the experiments, and the simulation environment is Pytorch 1.13.1 and NVIDIA GPU A100.
IV-A Neural Network Settings
As aforementioned that the DeepSC-ST 2.0 and the semantic-aware network are trained independently. In the DeepSC-ST 2.0, seven 1D CNN modules and 12 transformer encoder layers are concatenated to construct the semantic encoder. Three dense layers with 512 units are utilized in both the channel encoder and the channel decoder. The semantic decoder consists of a dense layer containing 32 units followed by a softmax activation function. In the semantic-aware network, three dense layers with 32 units are leveraged. The ChanEst network includes three 2D CNN layers containing 64 units, and the kernel size of each layer is $5\times 5$. The batch size is 64, and the number of warm-up steps is 3000 for training all the neural networks. The hyperparameter $\mu=0.3$. The parameter settings of the SAC-ST and the ChanEst network are summarized in Table I. Moreover, the trainable NN complexity of them is presented in Table II.
It is worth mentioning that the proposed SAC-ST is the first semantic MIMO system for speech transmission. For performance comparison, the DeepSC-ST [20] and the system developed in [18] are considered as two benchmarks. These two systems were initially designed for SISO channels, and we have refined them into MIMO transmission scenarios.
@@ IV-A Neural Network Settings
SAC-ST DeepSC-ST 2.0 Semantic Encoder 7$\times$1D CNN 512 channels GELU
12×12$\times$12 ×Transformer 128 (8 heads) GELU
Channel Encoder 2×2$\times$2 ×Dense Layer 512 units ReLU6
Dense Layer 512 units None
Channel Decoder 3×3$\times$3 ×Dense Layer 1024 units ReLU6
Semantic Decoder Dense Layer 32 units Softmax
Semanitc-Aware Network 2×2$\times$2 ×Dense Layer 32 units ReLU6
Dense Layer 32 units Softmax
ChanEst Network 2×2$\times$2 ×2D CNN 64 channels HardTanh
2D CNN 64 channels None
TABLE II: Neural network complexity of different systems.
Trainable Parameters
SAC-ST DeepSC-ST 2.0 $9.12\times 10^{7}$
Semanitc-Aware Network $3.17\times 10^{3}$
ChanEst Network $2.11\times 10^{5}$
@@ IV-B SU-MIMO Experiments
The WER results of the proposed SAC-ST in the SU-MIMO transmission system are shown in Fig. 7 (a), where perfect CSI is assumed. The ground truth is the WER score obtained by feeding the speech input into the speech-to-text pipeline, Conformer [47], directly and no speech transmission problems are considered. From the figure, the upgraded DeepSC-ST 2.0 achieves lower WER scores than the two benchmarks due to the improvement of the attention mechanism by incorporating the 1D CNN and transformer modules. The proposed SAC-ST outperforms the DeepSC-ST 2.0 and obtains a decrease of around $75.5\%$ when SNR$=$-8 dB. This is attributed to the semantic-aware network that transmits the essential semantic information over good SISO subchannels. Besides, the minimum SNR value to enable the readable sentence in the SAC-ST is around -8 dB, while it is around 4 dB in the DeepSC-ST 2.0. Moreover, the gap between the WER scores of the DeepSC-ST 2.0 and the proposed SAC-ST gradually narrows when the SNR value increases because the important semantic features in $\boldsymbol{f}$ are recovered accurately in both schemes when SNR$>$13 dB.
@@
Fig. 7 (b) illustrates the sentence similarity scores of different approaches. From the figure, the sentence similarity score of the DeepSC-ST and the DeepSC-ST 2.0 is nearly 0.2 for SNR$=$-8 dB, while the SAC-ST attains a score of 0.6. Similar to the result of WER, the proposed SAC-ST significantly outperforms the DeepSC-ST 2.0 in the low SNR regime, and the difference between them decreases when SNR$>$13 dB, which proves the effectiveness of the semantic-aware network in improving the semantic fidelity by perceiving the semantic importance over MIMO transmission paradigm with SVD precoding. The representative results of the DeepSC-ST, the work developed in [18], the DeepSC-ST 2.0, and the SAC-ST are shown in Table III, where the error words are highlighted in red. From the table, the sentence of the DeepSC-ST is corrupted and unable to read, and limited information is obtained from a few correct words scattered sporadically throughout the sentence of the DeepSC-ST 2.0. However, the sentence of the proposed SAC-ST is very intelligible and exhibits high semantic fidelity.
(a) WER
(b) Sentence similarity
@@
Next, we investigate the antenna diversity of SAC-ST to cope with various antenna distributions in realistic SU-MIMO communication scenarios, and the results of WER scores and sentence similarity scores are shown in Fig. 8. From the figure, SAC-ST has superior performance and is more robust to channel impairments in the transmission paradigm with more antennas. To verify that, we calculate the average equivalent SNR values, $\widetilde{\boldsymbol{\sigma}}$, in all SISO subchannels for different antenna combinations, and the example when true SNR value $\sigma$ is -8 dB is shown in Table IV [^1] All the data of the equivalent SNR values $\widetilde{\boldsymbol{\sigma}}$ for different antenna combinations and for true SNR value $\sigma$ from -8 dB to 22 dB is available at https://github.com/Zhenzi-Weng/SAC-ST/tree/main/equivalent-SNR.. From the Table, the SNR values in the top SISO subchannels of the antenna combination of $16\times 16$ are much higher than that of $2\times 2$, which facilitates more accurate semantic fidelity transmission. Moreover, as the number of antennas increases, the performance improvement between two different antenna combinations becomes negligible due to the insignificant difference between their SNR values in the top SISO subchannels.
TABLE III: Recovered sentences of the benchmarks, the DeepSC-ST 2.0, and the SAC-ST when SNR is -2 dB.
@@
Han’s work [18] rom th accepted cod of morcals in oden communities where tee doncant economic and egal featur of the menitysr lie s the instituton of private propery oneof the saia n features of the cd f moas is he sacedness of prperty
@@
TABLE IV: Equivalent SNR values $\widetilde{\boldsymbol{\sigma}}$ for different antenna combinations when true SNR value $\sigma$ is -8 dB.
@@
$N\times M$ Equivalent SNR values $\widetilde{\boldsymbol{\sigma}}$ in SISO subchannels (dB)
@@
The experimental results of the SAC-ST with different channel estimation algorithms are shown in Fig. 9. We consider three benchmarks. The first and second benchmarks test the SAC-ST trained under perfect CSI in the SU-MIMO transmission system with the conventional least square (LS) and minimum mean-square error (MMSE) channel estimation algorithms, respectively. The third benchmark is obtained by training and testing the SAC-ST under the channel environments estimated through the MMSE algorithm. From the figure, the SAC-ST with the ChanEst network outperforms all the benchmarks, especially in the low SNR regime, and it attains the WER scores and the sentence similarity scores comparable to the SAC-ST with perfect CSI. The representative results of the SAC-ST with various channel estimation methods in the SU-MIMO transmission system are shown in Table V.
TABLE V: Recovered sentences of the SAC-ST with different channel estimation algorithms when SNR is -8 dB.